Learners need to know the properties of numbers and the properties of operations on numbers. Several properties are important here:

• The commutative property of numbers for addition and multiplication.

• The associative property of numbers for addition and multiplication.

• The distributive property of numbers for multiplication over addition.

  • Commutative property Example: 4 p + 6 = 6 + 4 p

Do the following: 5 + 3 q = Example: 2 r × 3 s = 3 s × 2 r Do the following: 4 t × 6 =

• Associative property Example: (5 + 4 p) + (6 + 2 p) = (4 p + p 2 )+ (5 + 6)

Do the following: (4 q+ 5) + (3 q + 7) =

Example:

(5 r × 2 r) × ( 3 s × 4 s)

= (5 r × 3 s) × (2 r × 4 s) Do the following:

(4s×6)×(5×3s)=



1
Expert's answer
2021-05-25T17:45:13-0400

1. 5+3q=3q+5

2. 4t*6=6*4t

3. (4q+5)+(3q+7)=(5+7)+(4q+q 3)

4. (4s*6)*(5*3s)=(4s*5)*(6*3s)